OpenStudy (chealyn98):

A bicycle takes 8.0 seconds to accelerate at a constant rate from rest to a speed of 4.0 m/s. If the mass of the bicycle and rider together is 85 kg, what is the net force acting on the bicycle? (Hint: first calculate the acceleration.) Show your work

OpenStudy (anonymous):

This should probably be in the physics section not math but... anyway. Ok first try to identify relevant equations you might need here: I.E. Identify the variables you are given, the variables you need to solve for, and the mathematical relationships that connect them.

OpenStudy (chealyn98):

8.0 seconds 4.9 m/s 85 kg

OpenStudy (chealyn98):

A= 0.5 / 8 = 0.5 0.5 * 85 = 42.5 N